# RelianceScope: An Analytical Framework for Examining Students’ Reliance on Generative AI Chatbots in Problem Solving

### v1

@@ Abstract.
Reliance on AI, Analytical Framework, Chatbot-assisted Problem-Solving
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interaction design
Figure 1. RelianceScope is a framework for analyzing students’ chat and artifact edit logs in chatbot-assisted problem-solving activities through the lens of their knowledge contexts and reliance patterns. The nine reliance patterns are combinations of engagement modes in help-seeking and response-use, derived from the interaction model.
@@ 1. Introduction
Students’ cognitive engagement is particularly important for characterizing their reliance on a chatbot, given their interactions with it. We conceptualize student–chatbot interaction during problem-solving as a loop inspired by Norman’s execution-evaluation model (Norman, 2013): a student seeks help from a chatbot and then uses the chatbot’s response to modify a learning artifact. Within this loop, students can rely on chatbots at two key points. First, during help-seeking, students may rely on chatbots by delegating decisions about the type of help to receive (e.g., “help me” vs. “give me an example instead of the answer”) (Reeve and Tseng, 2011; Satyanarayan and Jones, 2024). Second, during response-use, students may rely on chatbots by delegating the completion of the learning task (e.g., copying a response vs. writing independently) (Chi and Wylie, 2014; Weatherholtz et al., 2025). Together, the engagement modes during help-seeking and response-use provide a holistic view of how students rely on chatbots throughout learning.
@@
We propose RelianceScope (Figure 1), an analytical framework for capturing and examining students’ joint reliance on chatbots during help-seeking and response-use. We defined three engagement modes (Chi and Wylie, 2014) for each, yielding 3 × 3 reliance patterns at individual student-chatbot interaction units. Besides, compared with prior approaches, RelianceScope further introduces a 2 × 2 knowledge-context lens that accounts for students’ mastery of the knowledge component and its instructional significance. The knowledge context enables more contextualized interpretations of reliance behaviors. Last but not least, analyses using RelianceScope are both scalable and analytically rich, as the framework relies on automatically collectable chat logs and artifact-edit logs without disrupting students’ learning processes.
@@
To the best of our knowledge, this work presents the first interaction-centric framework that integrates help-seeking and response-use to define reliance in generative chatbot-assisted learning. We envision that the fine-grained, context-sensitive, and non-intrusive design of RelianceScope will inform future investigations of reliance and guide the design of educational systems that better support self-regulated learners.
@@
  • A chatbot-assisted learning system deployed in a college-level web programming course that collects chat logs, code edit logs, and copy logs.
@@ 2.1. Reliance in Generative AI-Assisted Learning
Together, these mixed findings may suggest that the impact of AI reliance depends on contextual factors and interaction patterns that existing measures do not fully capture, underscoring the need for standardized and context-sensitive analytical tools (Ibrahim et al., 2025). RelianceScope addresses this gap by conceptualizing reliance along the dimensions of help-seeking and response-use—perspectives largely unexplored jointly in prior work—and by introducing a knowledge-context lens that enables systematic examination of when and under what conditions different reliance patterns emerge and affect learning.
@@ 2.2. Capturing Reliance on AI
Capturing how learners rely on AI is critical for understanding the role of chatbots in learning. Prior work has proposed taxonomies to characterize reliance from different perspectives. Many studies categorize reliance by examining the content and intent of learners’ messages (Graesser and Person, 1994; Yang et al., 2024; Sheese et al., 2024; Han et al., 2024; Lyu et al., 2024; McNichols et al., 2025), help-seeking behavior (Nelson-Le Gall, 1981; Karabenick, 2003; Kazemitabaar et al., 2023), or patterns of social interaction (Borchers et al., 2024). Others move beyond individual message turns to identify broader reliance profiles (Güner and Er, 2025; Stojanov et al., 2024; Hao et al., 2025).
@@ 3. RelianceScope
Contextual. We are particularly interested in capturing how students’ reliance on generative AI varies across learning contexts. For instance, learners may rely more on AI when limited prior knowledge constrains their ability to actively engage in an activity. Understanding how knowledge mastery and instructional significance shape reliance patterns can inform the design of adaptive generative-AI tutoring systems. Our framework leverages these knowledge contexts as a lens to enhance its analytical capabilities.
@@ 3.1.1. Reliance Patterns
The primary distinction between passive and active modes lies in the level of cognitive effort required to formulate requests and integrate responses (see Tables 1 and 2). The passive mode involves copying instructions and responses verbatim to ask questions and complete tasks, or waiting for next guidance. The active mode requires observable attention to part of knowledge, such as asking specific questions and critically analyzing responses. The constructive mode further involves learner agency, such as strategically shaping the form of assistance and independently creating new artifacts from AI responses. We exclude the interactive mode from the original ICAP framework, as it does not map well to chatbot-based, task-oriented learning interactions (Weatherholtz et al., 2025). By combining the three engagement modes in help-seeking and response-use, we derived nine distinct reliance patterns (Figure 1).
@@ 3.2. Classifying Reliance Patterns
Second, classifiers should consider adjacent interaction segments when determining the engagement mode. Engagement with a given knowledge component may be revealed in its subsequent segment rather than within itself. For instance, in Figure 2, the student initially discusses one knowledge component (“methods”) and then, in the subsequent segment, extends the conversation to a related component (“JavaScript”). Because the student builds on the initial explanation by asking follow-up questions, engagement with the first knowledge component is annotated as Active. In contrast, engagement with the second knowledge component is annotated as Passive, as the student reads the chatbot’s response without asking further questions or modifying the relevant part of the artifact.
@@ 4. Proof-of-Concept Study
We conducted a data-collection study in a college-level web programming course. Students in the course completed a programming activity using an assistive chatbot to learn Vue.js [^2] https://vuejs.org. We collected and classified students’ chat logs and code edit logs. To address RQ1, we conducted statistical analyses; for RQ2, we surveyed students; for RQ3, we benchmarked automatic classification.
@@ 4.1. Participants
The study involved 91 undergraduate students majoring in computer science, electrical engineering, and related fields. The students were taking an introductory web programming course in which they learn HTML, JavaScript, and Vue.js. A week before the study (academic week 11), students learned Vue.js in class. The students participated in the study as part of an assignment that affected their final grades. We collected students’ chat and code edit logs with informed consent and IRB approval. We excluded two students who declined to share data from the analysis.
@@ 4.2. Procedure
Before students interacted with a chatbot, we measured their prior knowledge (Stage 1 in Table 3) and self-regulation level (Stage 2). The pre-test consisted of 10 multiple-choice questions, each with a corresponding Vue.js knowledge component and an “I don’t know” option to minimize random guessing. The students were informed that their pre-test score would not affect their course grade, and they were encouraged to inform their unawareness to help us understand their prior knowledge. The pre-test had a 5-minute time limit, and the website displayed an alert prompting submission. The students then answered 12 questions about their self-regulation. We adopted the SLR-O questionnaire, specifically designed for online and hybrid courses (Broadbent et al., 2023). The students answered 3 indicative questions, each assessing self-efficacy, intrinsic motivation, extrinsic motivation, and metacognition, by rating their agreement with each statement on a 7-point Likert scale.
@@
After the learning activity, students took a post-test consisting of the same multiple-choice questions as the pre-test (Stage 5). The post-test also had a 5-minute time limit, and students were informed that their test score would impact their course grade. To further prevent random guessing, we informed students that we would deduct 1 point for incorrect answers, add 1 point for correct answers, and give 0 points for selecting “I don’t know.” At the end (Stage 6), students answered four free-form questions about their challenges in forming appropriate reliance on AI.
@@ 5.1. RQ1: Students’ Reliance Patterns
Figure 4. Flow of engagement modes from help-seeking to response use, with band widths representing proportions of interaction segments.
@@
Knowledge mastery did not alter reliance behaviors. We conducted a MANOVA to examine whether the distribution of the nine reliance patterns differed across three knowledge contexts: Acquired_Focal, Undeveloped_Focal, and Supporting. The independent variables were each student’s proportional use of the nine patterns within each context. Because these proportions are compositional, we applied a centered log-ratio (CLR) transformation with a multiplicative replacement strategy (Martín-Fernández et al., 2003) prior to analysis. The MANOVA revealed a significant effect of knowledge context (Pillai’s trace: $F=6.255,p<.001$). Follow-up ANOVAs with the Games-Howell post-hoc tests showed no significant differences between Acquired_Focal and Undeveloped_Focal, whereas both differed from Supporting for three reliance patterns as shown in Figure 5.
Similar reliance patterns across knowledge mastery are concerning because they suggest that students did not strategically allocate cognitive effort based on their learning needs. Instead of relying on the chatbot to handle acquired knowledge and focusing effort on undeveloped knowledge, students engaged similarly across both. This pattern may stem from three factors: students’ limited awareness of their own knowledge gaps, the scaffolding effect of prior knowledge that enables deeper engagement, or the chatbot’s uniform interaction regardless of students’ mastery levels. Future work should disentangle these factors to inform adaptive interventions that better support strategic reliance.
Figure 5. Distribution of reliance patterns across three knowledge contexts. Engagement modes on the x-axis are abbreviated as P (Passive), A (Active), and C (Constructive). The proportion of the P_P, P_A, and P_C patterns differ significantly between Focal and Supporting.
@@
The coefficient for Passive_Constructive was larger than that for Active_Passive, suggesting that deeper engagement during response-use may have a stronger association with learning than active help-seeking alone. However, because other coefficients did not reach statistical significance, we refrain from drawing definitive conclusions about the relative effectiveness of all reliance patterns.
Table 4. Ordinary least squares regression coefficients for the nine reliance patterns (formatted as {help-seeking mode}-{response-use mode}), predicting students’ post-test score. Asterisks denote statistically significant coefficients ($p<.05$).
@@
Figure 6. Transitions between reliance patterns. Engagement modes are abbreviated as P (Passive), A (Active), and C (Constructive).
@@ 5.2. RQ2: Challenges in Managing Reliance
Students struggled to pinpoint knowledge gaps and shape the chatbot’s behavior for constructive help-seeking. 17 of 91 students reported difficulty determining what to ask and how to ask. Some struggled to isolate the specific knowledge gaps within complex implementation steps that involved multiple knowledge components. As a result, they asked broad or underspecified questions, which led the chatbot to provide more hints or complete solutions than intended. Several students wanted to pose “questions more geared toward learning rather than finding answers (S7),” but faced “wording challenges” to prevent the chatbot from revealing answer code (S79). Twelve students struggled to align the chatbot with their personalized learning goals. Since students were using the chatbot for the first time, they were uncertain about its capabilities and how to “write prompts to maximise its value (S86).” The uncertainty limited their constructive help-seeking, such as steering the chatbot to provide brief answers to “see what was wrong and what needed to be fixed and why (S58)” or to provide more assistance with unfamiliar Vue syntax (S18).
@@ 5.3. RQ3: Scalability of RelianceScope
Few-shot examples improve LLM classification performance. Increasing the number of few-shot examples (from zero-shot to 3-shot and 9-shot) consistently improves classification performance across the five models, except Gemini-3-pro. This trend suggests that performance can be further enhanced as additional annotated data becomes available through our data collection system. Notably, reasonable performance is achieved with as few as nine shots, indicating that automatic reliance classification is both generalizable and accessible, without requiring large-scale data collection or activity-specific fine-training. Performance may be further improved by incorporating lightweight rule-based signals for certain patterns. For example, high textual similarity between a student’s message and the activity instructions may indicate Passive help-seeking, and the presence of copy logs immediately following a chatbot response may signal Passive response-use.
@@ 6. Discussion
We discuss how RelianceScope supports a deeper understanding of AI reliance and provide design guidelines for learning systems.
@@ 6.1. Implications of RelianceScope
The significance of RelianceScope lies in its interaction-centric design, which jointly models help-seeking and response-use to capture reliance holistically. Prior work has largely examined these behaviors in isolation—an approach suited to structured tutoring systems that constrain either how students ask for help or how they consume feedback, but ill-suited to the open-ended nature of generative AI chatbots. Our analysis shows that 46% of interactions exhibit different engagement modes in help-seeking and response-use, including sharply contrasting patterns (e.g., Passive_Constructive). For example, one student (S37) repeatedly engaged in passive help-seeking by submitting task instructions to the chatbot, yet constructively used its responses to independently write code. Such hybrid reliance patterns fall between full reliance and full independence and would be obscured by existing measures, despite being critical for understanding how reliance manifests in practice. While prior work suggests that constructive response-use may compensate for passive help-seeking (Aleven et al., 2016) and that guided learning can outperform self-exploration (Yannier et al., 2020), RelianceScope provides a systematic means to investigate the interplay between help-seeking and response-use more thoroughly.
@@ 6.2. Design Implications for Learning Systems
We share several directions for designing future AI-assisted learning systems. Most college-level students in the survey reported awareness of the risks of over-reliance on AI and an intention to self-regulate their use of chatbots. However, students lacked adequate support for decomposing learning tasks and requesting assistance only for the specific components they had not yet mastered—capabilities that are critical for incremental learning. These gaps indicate opportunities for intervention at both the learning-activity and chatbot-design levels.
At the level of learning activities, we note the value of incremental task design. In our study, the programming assignment was decomposed into six sub-activities, revealed sequentially. Notably, none of the 79 students asked the chatbot to generate the entire solution at once. This incremental structure likely helped students more accurately assess their knowledge state and created more opportunities to adjust their help-seeking and reliance on AI over time. Such task-level scaffolding may be more robust than chatbot-only interventions, as students can easily opt out of institutional chatbots in favor of general-purpose tools, whereas they must engage with the structure of the assigned activity.

### v2

@@ Abstract.
Reliance on AI, Analytical Framework, Problem-Solving
journalyear: 2026copyright: ccconference: Proceedings of the Thirteenth ACM Conference on Learning @ Scale; June 29-July 03, 2026; Seoul, Republic of Koreabooktitle: Proceedings of the Thirteenth ACM Conference on Learning @ Scale (L@S ’26), June 29-July 03, 2026, Seoul, Republic of Koreadoi: 10.1145/3774398.3811612isbn: 979-8-4007-2293-6/2026/06ccs: Human-centered computing Interaction design
Figure 1. Combined engagement in help-seeking and response-use is critical for examining students’ reliance on AI during chatbot-assisted problem-solving. RelianceScope is a framework for analyzing these reliance patterns from chat and artifact edit logs through the lens of knowledge contexts.
@@ 1. Introduction
Students’ cognitive engagement is particularly important for characterizing their reliance on a chatbot, given their interactions with it. We conceptualize student–chatbot interaction during problem-solving as a loop inspired by Norman’s execution-evaluation model (Norman, 2013): a student seeks help from a chatbot and then uses the chatbot’s response to modify a learning artifact. Within this loop, students can rely on chatbots at two key points. First, during help-seeking, students may rely on chatbots by delegating decisions about the type of help to receive (e.g., “help me” vs. “give me an example instead of the answer”) (Reeve and Tseng, 2011; Satyanarayan and Jones, 2024). Second, during response-use, students may rely on chatbots by delegating the completion of the learning task (e.g., copying a response vs. writing independently) (Chi and Wylie, 2014; Weatherholtz et al., 2025). Together, the engagement modes during help-seeking and response-use provide a holistic view of how students rely on chatbots.
@@
We propose RelianceScope (Figure 1), an analytical framework for capturing and examining students’ joint reliance on chatbots during help-seeking and response-use. We defined three engagement modes (Chi and Wylie, 2014) for each, yielding 3 × 3 reliance patterns at individual student-chatbot interaction units. Besides, compared with prior approaches, RelianceScope further introduces a 2 × 2 knowledge-context lens that accounts for students’ mastery of the knowledge component and its instructional significance. The knowledge context enables more contextualized interpretations of reliance behaviors. Last but not least, analyses using RelianceScope are both scalable and analytically rich, as the framework relies on automatically collectible chat logs and artifact-edit logs without disrupting students’ learning processes.
@@
To the best of our knowledge, this work presents the first interaction-centric framework that integrates help-seeking and response-use to define reliance in generative AI chatbot-assisted learning. We envision that the fine-grained, context-sensitive, and non-intrusive design of RelianceScope will inform future investigations of reliance and guide the design of educational systems that better support self-regulated learners.
@@
  • A generative AI chatbot-assisted learning system deployed in a college-level web programming course that collects chat logs, code edit logs, and copy logs.
@@ 2.1. Reliance in Generative AI-Assisted Learning
Together, these mixed findings may suggest that the impact of AI reliance depends on contextual factors and interaction patterns that existing measures do not fully capture, underscoring the need for standardized and context-sensitive analytical tools (Ibrahim et al., 2025). RelianceScope addresses this gap by conceptualizing reliance along the dimensions of help-seeking and response-use—perspectives largely unexplored jointly in prior work—and by introducing a knowledge-context lens that enables systematic examination of when different reliance patterns emerge and affect learning.
@@ 2.2. Capturing Reliance on AI
Capturing how learners rely on AI is critical for understanding the role of chatbots. Prior work has proposed taxonomies to characterize reliance from different perspectives. Many studies categorize reliance by examining the content and intent of learners’ messages (Graesser and Person, 1994; Yang et al., 2024; Sheese et al., 2024; Han et al., 2024; Lyu et al., 2024; McNichols et al., 2025), help-seeking behavior (Nelson-Le Gall, 1981; Karabenick, 2003; Kazemitabaar et al., 2023; Myung et al., 2026), or patterns of social interaction (Borchers et al., 2024). Others move beyond individual message turns to identify broader reliance profiles (Güner and Er, 2025; Stojanov et al., 2024; Hao et al., 2025).
@@ 3. RelianceScope
Context-sensitive. We are particularly interested in capturing how students’ reliance on generative AI varies across learning contexts. For instance, learners may rely more on AI when limited prior knowledge constrains their ability to actively engage in an activity. Understanding how knowledge mastery and instructional significance shape reliance patterns can inform the design of adaptive generative-AI tutoring systems. Our framework leverages these knowledge contexts as a lens to enhance its analytical capabilities.
@@ 3.1.1. Reliance Patterns
The primary distinction between passive and active modes lies in the level of cognitive effort required to formulate requests and integrate responses (see Tables 1 and 2). The passive mode involves copying instructions and responses verbatim to ask questions and complete tasks, or waiting for the next guidance. The active mode requires observable attention to part of knowledge, such as asking specific questions and critically analyzing responses. The constructive mode further involves learner agency, such as strategically shaping the form of assistance and independently creating new artifacts from AI responses. We exclude the interactive mode from the original ICAP framework, as it does not map well to chatbot-based, task-oriented learning interactions (Weatherholtz et al., 2025). By combining the three engagement modes in help-seeking and response-use, we derived nine distinct reliance patterns (Figure 1).
@@ 3.2. Classifying Reliance Patterns
Second, classifiers should consider adjacent interaction segments when determining the engagement mode. Engagement with a given knowledge component may be revealed in its subsequent segment rather than within itself. For instance, in Figure 2, the student initially discusses one knowledge component (“methods”) and then, in the subsequent segment, extends the conversation to a related component (“JavaScript”). Because the student builds on the initial explanation by asking follow-up questions, the first interaction segment is annotated as Active response-use. In contrast, the second is annotated as Passive response-use, as the student neither asked further questions nor modified the relevant part of the artifact.
@@ 4. Proof-of-Concept Study
We conducted a data-collection study in a college-level web programming course. Students in the course completed a programming activity using an assistive chatbot to learn Vue.js. We collected and classified students’ chat logs and code edit logs. To address RQ1, we conducted statistical analyses; for RQ2, we surveyed students; for RQ3, we benchmarked automatic classification.
@@ 4.1. Participants
The study involved 91 undergraduate students majoring in computer science, electrical engineering, and related fields. The students were taking an introductory web programming course in which they learned HTML, JavaScript, and Vue.js. A week before the study (academic week 11), students learned Vue.js in class. The students participated in the study as part of an assignment that affected their final grades. We collected students’ chat and code edit logs with informed consent and IRB approval. We excluded two students who declined to share data from the analysis.
@@ 4.2. Procedure
Before students interacted with a chatbot, we measured their prior knowledge (Stage 1 in Table 3). The pre-test consisted of 10 multiple-choice questions, each with a corresponding Vue.js knowledge component and an “I don’t know” option to minimize random guessing. The students were informed that their pre-test score would not affect their course grade. The pre-test had a 5-minute time limit, and the website displayed an alert prompting submission. The students then reported their self-efficacy, intrinsic motivation, extrinsic motivation, and metacognition (Stage 2) by answering 12 7-point Likert-scale questions (Broadbent et al., 2023).
@@
After the learning activity, students took a post-test consisting of the same multiple-choice questions as the pre-test (Stage 5). The post-test also had a 5-minute time limit, and students were informed that their test score would impact their course grade. To further prevent random guessing, we informed students that we would deduct 1 point for incorrect answers. At the end (Stage 6), students answered four free-form questions about their challenges in forming appropriate reliance on AI.
@@ 5.1. RQ1: Students’ Reliance Patterns
Figure 4. Flow of engagement modes from help-seeking to response-use, with band widths representing proportions of interaction segments.
@@
Knowledge mastery did not alter reliance behaviors. We conducted a MANOVA to examine whether the distribution of the nine reliance patterns differed across three knowledge contexts: Acquired_Focal, Undeveloped_Focal, and Supporting. The independent variables were each student’s proportional use of the nine patterns within each context. Because these proportions are compositional, we applied a centered log-ratio (CLR) transformation with a multiplicative replacement strategy (Martín-Fernández et al., 2003) prior to analysis. The MANOVA revealed a significant effect of knowledge context (Pillai’s trace: $F=6.255,p<.001$). Follow-up ANOVAs with Games-Howell post-hoc tests showed no significant differences between Acquired_Focal and Undeveloped_Focal, whereas both differed from Supporting in three reliance patterns (Figure 5).
Similar reliance patterns across knowledge mastery are concerning because they suggest that students did not strategically allocate cognitive effort. Students could have relied more on the chatbot for Acquired knowledge and devoted more effort to practicing Undeveloped knowledge. This pattern may reflect students’ limited awareness of knowledge gaps, insufficient metacognition to select effective strategies, or the chatbot’s uniform interactions across mastery levels.
Figure 5. Distribution of reliance patterns across three knowledge contexts. Engagement modes are labeled by their first letter (e.g., P for Passive). The proportions of P_P, P_A, and P_C differ significantly between Focal and Supporting.
@@
The coefficient for Passive_Constructive (0.371) was more positive than that for Active_Passive (-0.615), suggesting that deeper engagement during response-use may have a stronger association with learning than active help-seeking alone. However, because other coefficients did not reach statistical significance, we refrain from drawing definitive conclusions about the relative effectiveness of all reliance patterns.
Table 4. Ordinary least squares regression coefficients for the nine reliance patterns, predicting students’ post-test score. Asterisks denote statistically significant coefficients ($p<.05$).
@@
Figure 6. Red arrows show more common transitions between reliance patterns; blue arrows show less common ones.
@@ 5.2. RQ2: Challenges in Managing Reliance
Students struggled to pinpoint knowledge gaps and shape the chatbot’s behavior for constructive help-seeking. 17 of 91 students reported difficulty determining what to ask and how to ask. Some struggled to isolate the specific knowledge gaps within complex implementation steps that involved multiple knowledge components. As a result, they asked broad or underspecified questions, which led the chatbot to provide more hints or complete solutions than intended. Several students wanted to pose “questions more geared toward learning rather than finding answers (S7),” but faced “wording challenges” to prevent the chatbot from revealing answer code (S79). Twelve students struggled to align the chatbot with their personalized learning goals. Since students were using the chatbot for the first time, they were uncertain about its capabilities and how to “write prompts to maximize its value (S86).” The uncertainty limited their constructive help-seeking, such as steering the chatbot to provide brief answers to “see what was wrong and what needed to be fixed and why (S58)” or to provide more assistance with unfamiliar Vue syntax (S18).
@@ 5.3. RQ3: Scalability of RelianceScope
Few-shot examples improve LLM classification performance. Increasing the number of few-shot examples (from zero-shot to 3-shot and 9-shot) consistently improves classification performance across the five models, except Gemini-3-pro. This trend suggests that performance can be further enhanced as additional annotated data becomes available through our data collection system. Notably, reasonable performance is achieved with as few as nine shots, indicating that automatic reliance classification is both generalizable and accessible, without requiring large-scale data collection or activity-specific fine-tuning. Performance may be further improved by incorporating lightweight rule-based signals for certain patterns. For example, high textual similarity between a student’s message and the activity instructions may indicate Passive help-seeking, and the presence of copy logs immediately following a chatbot response may signal Passive response-use.
@@ 6.1. Implications of RelianceScope
The significance of RelianceScope lies in its interaction-centric design, which jointly models help-seeking and response-use to capture reliance holistically. Prior work has largely examined these behaviors in isolation—an approach suited to structured tutoring systems that constrain either how students ask for help or how they consume feedback, but ill-suited to the open-ended nature of generative AI chatbots. Our analysis shows that 46% of interactions exhibit different engagement modes in help-seeking and response-use, including sharply contrasting patterns (e.g., Passive_Constructive). For example, one student (S37) repeatedly engaged in passive help-seeking by submitting task instructions to the chatbot, yet constructively used its responses to independently write code. Such hybrid reliance patterns fall between full reliance and full independence and would be obscured by existing measures, despite being critical for understanding how reliance manifests in practice. While prior work suggests that constructive response-use may compensate for passive help-seeking (Aleven et al., 2016) and that guided learning can outperform self-exploration (Yannier et al., 2020), RelianceScope provides a systematic means to investigate the interplay between help-seeking and response-use.
@@ 6.2. Design Implications for Learning Systems
We share several directions for designing future AI-assisted learning systems. Most college-level students in the survey reported awareness of the risks of over-reliance on AI and an intention to self-regulate their use of chatbots. However, students lacked adequate support for decomposing learning tasks and requesting assistance only for the specific components they had not yet mastered. These gaps indicate opportunities for intervention at both the learning-activity and chatbot-design levels.
At the level of learning activities, we note the value of incremental task design. In our study, the programming assignment was decomposed into six sub-activities, revealed sequentially. Notably, none of the 79 students asked the chatbot to generate the entire solution at once. This incremental structure likely helped students more accurately assess their knowledge state and created more opportunities to adjust their help-seeking and reliance on AI over time. Such task-level scaffolding may be more robust than improving chatbots only, as students can easily opt out of institutional chatbots in favor of general-purpose tools, whereas they cannot bypass incrementally revealed instructions.